OpenStudy (anonymous):

2(4x+3)=3(2x-4) helpp. show steps.

OpenStudy (anonymous):

Open the parentheses: 8x + 6 = 6x - 12 Move all the variables to the same side: 8x - 6x = -12 - 6 2x = - 18 x = -9